Comprehensive Problem (Chapter 2) Extreme Edge Ine. Income Statements For the Year Ended December 31 Sales revenues 2015 2014 S 1,310 1,240 Cost of goods sold Gross profit Depreciation expense 510 136 369 521 120 341 Income from continuing operations Financial (interest) expense Income tax expense (benefit) Net income (loss) 24 0.48 Earnings (loss) per share Statements of Shareholders Equity (in thousands, except par value) Balance as of January 1, 2014 Common stock issued, $1 par value Net income Dividends declared and paid Common Additional Retained Stock PIC 50 S 450 S 24 2) Balance as of December 31, 2014 50 S 450 S 50 450 S Balance as of January 1, 2015 Common stock issued, S1 par value Net income (loss) Dividends declared and paid Balance as of December 31. 2015 Balance Sheets at December 31 50 S 450 2015 2014 Current Assets: Accounts receivable, net Inventory Prepaid expenses 160 335 187 175 280 Total current assets 655 Property, Plant and Equipment: Equipment, net of depreciation 204- 260 915 Total Assets S 806Current Liabilities: Accounts payable 110 Accrued liabilities Total current liabilities Long-term Liabilities 49 149 129 Notes payable Total Liabilities Shareholders Equity Common stock, $1 par Additional paid-in-capital on common stock Total contributed capita Retained carnings Total Shareholders Equity 249 450 450 Total Liabilities and Shareholders Equity 915 Statements of Cash Flows For the Year Ended December 31 Cash flows from operating activities 2015 2014 (9) S 136 Net Income Depreciation expense, equipment Accounts Receivable 120 Changes in current accounts: 15 (55) 30 (35) (15) Prepaid Expenses Accounts Payable Accrued Liabilities Net cash provided (used) by operating activitics 112 144 Cash flows from investing activities Purchase of equipment Net cash provided (used) by investing activities 140) Cash flows from financing activities 20) Issue (retire) notes payable Issue common stock Pay cash dividend Net cash provided (used) by financing activities Net change in cash (120) S (88) S 142 S 187 Cash, beginning of the year Net change in eash Cash, end of the year 88)
Answer #1

The answers are highlighted in bold below and their computations are also provided:

(1): Cost of goods sold in 2015 = sales revenue – gross profit = 1310-510 = 800

(2): Earnings (loss) per share in 2015 = net loss of 2015/no. of shares = -9/50 = -0.18

(3): Retained earnings balance as of December 31, 2105 = opening balance – net loss = 66-9 = 57

(4): Cash in 2015 = Total current assets – accounts receivable – inventory - prepaid expenses = 602-160-335 - 8 = 99

(5): Total liabilities and shareholders equity in 2015 = total liabilities+total equity = 249+557 = 806

(6): Cash end of the year = beginning cash-net change in cash = 187-88 = 99